OpenStudy (anonymous):

Input the equation of the given line in standard form. The line with m = -1/2 and the point (-2, -3).

OpenStudy (sw050399):

Remember the point slope form equation, \[y - y _{1} = m(x - x _{1})\]?

OpenStudy (anonymous):

m = -1/2 (x1,y1) = (-2,-3) y-y1 = m(x-x1) plug the value of x1, y1 and m in the above equation and then write it in the from Ax +By = C

OpenStudy (anonymous):

y=1/2x-4

OpenStudy (anonymous):

y = -1/2x -4

OpenStudy (anonymous):

this is the slope intercept form write this equation in standard form

OpenStudy (anonymous):

1/2x+y=-4

OpenStudy (anonymous):

yes its right now multiply both sides by 2 to get rid of 1/2

OpenStudy (anonymous):

x+2y=-8

OpenStudy (anonymous):

yes its the required standard from

OpenStudy (anonymous):

*form

OpenStudy (anonymous):

can you help me with another?

OpenStudy (anonymous):

yes sure

OpenStudy (anonymous):

Input the equation of the given line in standard form. The line including the two points (5, 3), (8, -2)

OpenStudy (anonymous):

midpoint is -5/3 ?

OpenStudy (anonymous):

no dont try to find the midpoint you just need the equation in standard form

OpenStudy (anonymous):

here is trick the slope is not given can you find the slope from the given two points???????

OpenStudy (anonymous):

????????????????????????????????????????????

OpenStudy (anonymous):

I don't know how...

OpenStudy (anonymous):

is the answer 5x+y=16

OpenStudy (anonymous):

you have these two points. (5, 3), (8, -2) use the following formula to calculate the slope m = (y2-y1)/(x2-x1) here y2 = -2 y1 = 3 x2 = 8 and x1 = 5

OpenStudy (anonymous):

dont go for the answer directly

OpenStudy (anonymous):

when you calculate the slope, you get =5/3 right/

OpenStudy (anonymous):

slope = -5/3

OpenStudy (anonymous):

now take one of the two points and find the equation by using the formula y-y1 = m(x-x1)

OpenStudy (anonymous):

y- -3=-5/3(x-5) y+3=-5/3x+25/3

OpenStudy (anonymous):

then you'd multiply by 3 on each side right?

OpenStudy (anonymous):

y-3 = -5/3(x-5) now simplify this equation

OpenStudy (anonymous):

y+3=-5/3x+25/3

OpenStudy (anonymous):

its not y+3 its y-3

OpenStudy (anonymous):

the rest is ok

OpenStudy (anonymous):

y-3=-5/3x+25/3 y +5/3x = 25/3 +3 multiply both sides by 3

OpenStudy (anonymous):

3y+5x=28 ?

OpenStudy (anonymous):

3y + 5x = 25 + 9 3y + 5x = 34 or 5x +3y = 34 -----> take your answer